# Chen, Yan, and Jiang

### v1

@@ 5 Conclusion
Motivated by growing pricing applications in centralized matching markets, we study the behavior of the value function of the linear program (LP) introduced by Aouad and Sarıtaç (2022) for cost-minimizing dynamic stochastic matching, with a particular focus on its (weak) concavity with respect to agents’ arrival rates. We establish a key result: it is sufficient for all optimal solutions of the LP to have strictly positive unmatched rates to ensure weak concavity. Leveraging this insight, we demonstrate that weak concavity is guaranteed when there are at most three agent types sharing the same patience level. For the general case with an arbitrary number of types, (weak) concavity is more likely to hold when agents are either highly patient, have similar patience levels, or exhibit substantial heterogeneity in matching efficiencies.

### v2

@@ 5 Conclusion
Motivated by growing pricing applications in centralized matching markets, we study the behavior of the value function of the LP introduced by Aouad and Sarıtaç (2022) for cost-minimizing dynamic stochastic matching, with a particular focus on its (weak) concavity with respect to agents’ arrival rates. We establish a key result: it is sufficient for all optimal solutions of the LP to have strictly positive unmatched rates to ensure weak concavity. Leveraging this insight, we demonstrate that weak concavity is guaranteed when there are at most three agent types sharing the same patience level. For the general case with an arbitrary number of types, (weak) concavity is more likely to hold when agents are either highly patient, have similar patience levels, or exhibit substantial heterogeneity in matching efficiencies.